OpenStudy (anonymous):

solve: square root of x+5 = square root of (x) +1

OpenStudy (tkhunny):

This? \(\sqrt{x+5} = \sqrt{x} + 1\) ABSOLUTELY, POSITIVELY the first thing to do is consider the Domain. \(\sqrt{x}\) insists that \(x \ge 0\) \(\sqrt{x+5}\) insists that \(x \ge -5\) After WRITING DOWN \(x \ge 0\), since one radical is already isolated, simply square both sides.

OpenStudy (anonymous):

thank you so much!